The transfer function relating Z(s) to the input U(s) is given below: It can be shown that the ILT of Z (s) is of the form Z(t) Find C1, C2, r1 r2 r3 , and WProblem 1: (5Points) The transfer function relating Z(s) to the input U(s) is given below: St S It can be shown that the ILT
Answer #1

Given,

Z(s)=\frac{s+5}{s^{2}+9s+20}U(s)

and

U(s)=\frac{9}{s^2+9}

so,

Z(s)=\frac{s+5}{s^{2}+9s+20}*\frac{9}{s^2+9}

and laplace inverse of Z(s) can be calculated as below

z(t)=L^{-1}\left\{\frac{9\left(s+5\right)}{\left(s^2+9s+20\right)\left(s^2+9\right)}\right\}

using partial fractions and converting into simpler forms

z(t)=L^{-1}\left\{\frac{36-9s}{25\left(s^2+9\right)}+\frac{9}{25\left(s+4\right)}\right\}

z(t)=L^{-1}\left\{-\frac{9s}{25\left(s^2+9\right)}+\frac{36}{25\left(s^2+9\right)}+\frac{9}{25\left(s+4\right)}\right\}

z(t)=\frac{12}{25}\sin \left(3t\right)-\frac{9}{25}\cos \left(3t\right)+\frac{9}{25}e^{-4t}